OpenStudy (anonymous):

Which of the following circles have their centers in the third quadrant? A. (x + 14)2 + (y - 14)2 = 84 B. (x + 16)2 + (y + 3)2 = 17 C. (x + 3)2 + (y - 6)2 = 44 D. (x + 9)2 + (y + 12)2 = 36

OpenStudy (anonymous):

The third quadrant has both a negative x and y coordinate. To find this, find the ones with both x+a number and y+a number since the form is (x−h)2+(y−k2)=r2 So im comfused ...

OpenStudy (triciaal):



OpenStudy (triciaal):

equation of a circle (x-h)^2 + (y-k)^2 = r^2 center (h,k) and radius r

OpenStudy (anonymous):

Thank you so much I finally get it now @triciaal

OpenStudy (triciaal):

great!!